Include your work by showing the steps you took to solve it or explain in words what you did for the problem.
Solve the following problem:
hint( If consecutive integers are defined as n,n+1, etc. how could you find EVEN integers be defined)
Problem:
The sum of four consecutive EVEN integers is 308. write and solve an equation to find four integer's.

note; im in cyber school and this is the third time ive got this problem wrong. and the first time i turned to this site. their answer was wrong. Can someone please help me.
Thankyou,
Ashley

Answer by sudhanshu_kmr(1152)


Let assume even integers are x,x+2, x+4, x+6
sum of these four = 308
=> x + x+2 +x+4 + x+6 = 308
=> 4x + 12 = 308
=> 4x = 308-12
=> x = 296/4 = 74
so, even integers are 74, 76, 78 and 80.